OpenStudy (anonymous):

(16 + 32 ÷ 4)(21 – 6 • 3)

OpenStudy (unanimoose):

Are you familiar with PEMDAS?

OpenStudy (anonymous):

nah

OpenStudy (mikeymaximum):

The 1st part is 24.75

OpenStudy (unanimoose):

OK, well. P = Parenthesis E = Exponent M/D = Multiplication/Division A/S = Addition/Subtraction

OpenStudy (mikeymaximum):

and the 2nd part is...

OpenStudy (unanimoose):

@MikeyMaximum Please don't be rude. I am already trying to help him and you come here blurting the answer.

OpenStudy (mikeymaximum):

...3

OpenStudy (unanimoose):

And you didn't even do it right. >_>

OpenStudy (anonymous):

i give u a trophe

OpenStudy (mikeymaximum):

I did the 1st () 1st and then I did the ones on the right

OpenStudy (mikeymaximum):

How is that doing it wrong?

OpenStudy (unanimoose):

You didn't do the first one right lol.

OpenStudy (unanimoose):

But since you clearly don't care about giving our answers (which is against the CoC), I'm just gonna go.

OpenStudy (anonymous):

Out of my question Mikey all i just want is Help

OpenStudy (unanimoose):

out answers*

OpenStudy (mikeymaximum):

U r supposed 2 do the division 1st

OpenStudy (unanimoose):

But how did you even get 24.75?

OpenStudy (mikeymaximum):

32/4=8

OpenStudy (mikeymaximum):

and...

OpenStudy (mikeymaximum):

nvm...

OpenStudy (unanimoose):

You know what, I can't do anything here with you doing all the work for him wrong. Bye.

OpenStudy (anonymous):

I GOT 72?

OpenStudy (mikeymaximum):

I WAS NOT TRYING 2 BE RUDE IN ANYWAY MOOSE btw U have earned a medal 4 finding mistakes XD Sowwy

OpenStudy (anonymous):

...

OpenStudy (mikeymaximum):

CORRECTION The 1st part is actually 24

OpenStudy (anonymous):

ok

OpenStudy (mikeymaximum):

And... can u write the 2nd part again becuz on my screen I have a WHOLE bunch of ?s